From nch@roe.ac.uk Fri Aug 13 16:39:50 2010 Date: Fri, 13 Aug 2010 16:03:43 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren , Tom Kerr Subject: WFAU VDFS Science Archives *fortnightly* meet mins, 13/08/10 Minutes of WFAU VDFS Science Archive meeting: August 13th 2010 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NJC, MAR, NCH, RSC, ETWS, KTN Apologies: JPE, AL, MSH, RGM, RPB ** NB: DoNM: Friday August 27th in the Vista Hut ** Actions discharged this week: ----------------------------- ACTION: ETWS to run his auto fitskeys check on the current schema to make sure no vital metadata is missing from the VSA Discharged; 6 new keys have appeared and have been added into the VSA template schema by MAR and ETWS. ACTION: NCH to check GPS DR6 for the odd missing merged frame and make a note on the WSA website as appropriate. Discharged; turns out there were a few missing frames so appropriate notes have been added to the web pages. ACTION: NCH to do his usual quick science validation/sanity checks on LAS/GCS DR8. Discharged; all looks fine Actions partly discharged but continuing: ----------------------------------------- ACTION: RSC & NJC to document new operational procedures on the TWiki Continues; further additions this week from RSC and NJC noted that these notes will be finalised once the DR8 UDS has been run through with the required changes. ACTION: NCH to assign actions on the following next time: - Modify entire code base to work with Tiles (Partly discharged - but needs fully testing on the new V1 data). This includes checking code works here on deep stacks, checking quality control/ QBF, ingest, CU7, CU6. Also do we need to do better links between tile detections and pawprint detections or will Provenance etc be good enough. (RSC has added tile qbits; also suggested a preliminary test run of CU7 using shallow tiles first of all). - Change schema / CU4 for new aperture mag names and columns for corrections (partly discharged by ETWS and to be finalised by NJC) - VVV SynopticSource (allowing both SourceXSynopticSourceBestMatch and SourceXDetectionBestMarch) - Updating Filter table with A/E(B-V) and VegaToAB (Need values from STH) Continues Actions carried forward from 06/08/10 meeting --------------------------------------------- ACTION: TBD to draft and MOU on behalf of PSPIs for ESO-SAF deliveries Continues; on hold until at least next PSPI teleconf, and possibly until "Phase 3" workshop at ESO. ACTION: NJC to suggest workshop at next PSPI meeting. ACTION: RPB to update WFCAMOPENTIME for 09A projects no longer inside their proprietary period. Specific points and new actions: -------------------------------- Project management: NCH suggested that from here onwards these meetings be convened on a fortnightly basis owing to increasing pressures from Gaia work. MAR in particular agreed with great enthusiasm, despite not working on Gaia (lucky fellow that he is). WFCAM & VISTA updates: Luca Rizzi steps down from his role as UKIDSS support astronomer at the JAC this week. We all wish him well in his new role in the JCMT software group, and thank him for his hard work in ensuring the success of the WFCAM surveys. In bocca al lupo! (hopefully the Internet is not misleading me with some inflamatory translation...) Comments and issues arising from CASU minutes: No new minutes this week. Networking: Transfers from CASU are all up-to-date: VISTA v1.0 up to and including May 24th; WFCAM up to and including all of May. WSA/VSA Operations: ETWS noted -- Re-transferred and re-ingested the data for 2009-07-29 where a detector was missing originally for some data. -- DR8 release progress: GCS and LAS (re-run) are finished and waiting for release. DXS and UDS are still running. -- VISTA data version 1.0 up to 2010-05-24 is transferred and metadata is ingested. Processing and ingest of P85 catalogue data has started, SV and DRYRUN catalogue data is ingested. -- Checked for new FITS keywords and updated schemas accordingly. NCH noted several email communications with Phil Lucas concerning a small fraction of fields in the GPS where the photometric calibration has gone wonky owing to large and variable extinction. A form of words for the web site has been provided, and also a new ppErrBit defined for flagging those catalogue detections as likely affected. ACTION: MAR to add some web documentation for the GPS photometric calibration issues. ACTION: RSC to add new ppErrBit to GPS catalogues as per instructions from the survey head; and deprecate affected framesets to ensure correct propagation/reseaming in gpsSource DR7. Following DR8, GPS DR7 will be done once the new quality error bits have been set. MAR noted that he'll start the VISTA auto-QC next week once the catalogue ingests are up-to-date. Hardware and Systems: MAR noted that NAS box takelot is a bit poorly, with a dodgy network card. Apparently as long as the power stays on it will continue to serve it's data (legacy supercos plus more recent WFCAM data) but may go permanently offline if inadvertently powered-down. (Eclipse computing have been called in and were supposed to fix on Wednesday this week, but cancelled at the last minute.) Software: RSC noted: "Implemented deep WFCAM stack catalogue saturation flagging using the correction criterion as applied to VISTA. The changes are documented on the website. Also, tested and fixed a few regression bugs in NJC's latest developments for automated programme set up of the DXS and deep curation changes to support the latest UDS mosaic." Concerning the memory leak identified by RSC in PyFITS 2.x, ETWS installed the release fix and RSC tested it; all looks OK now. Survey Data Release: Nothing further to add over notes above. Non-survey Data Release: Nothing of note this week. Astrogrid deployment & Data Analysis services: Nothing to report this week. Miscellaneous: NJC noted he has registered for ADASS this week. KTN noted that MSH may also be going and presenting a paper on grid-related things; RSC is still considering going. NCH noted that he's making a contribution to the PanSTARRS science workshop at QUB in a couple of weeks concerning the VDFS experience of crossmatching bulk datasets. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================